# Training Consistency Models with Variational Noise Coupling

### v1

@@ Abstract
Consistency Training (CT) has recently emerged as a promising alternative to diffusion models, achieving competitive performance in image generation tasks. However, non-distillation consistency training often suffers from high variance and instability, and analyzing and improving its training dynamics is an active area of research. In this work, we propose a novel CT training approach based on the Flow Matching framework. Our main contribution is a trained noise-coupling scheme inspired by the architecture of Variational Autoencoders (VAE). By training a data-dependent noise emission model implemented as an encoder architecture, our method can indirectly learn the geometry of the noise-to-data mapping, which is instead fixed by the choice of the forward process in classical CT. Empirical results across diverse image datasets show significant generative improvements, with our model outperforming baselines and achieving the state-of-the-art (SoTA) non-distillation CT FID on CIFAR-10, and attaining FID on par with SoTA on ImageNet at $64\times 64$ resolution in 2-step generation. Our code is available at https://github.com/sony/vct.
@@
Figure 1: Comparison of 1-step generation on toy data for independent and variational coupling. The data is sampled from a 2-d mixture of Gaussians with means $\bm{\mu}_{1}=(0,0.5)$ and $\bm{\mu}_{2}=(0,-0.5)$. On the bottom plot (our approach), we show the posterior probabilities learned by the encoder (in blue and green) corresponding to $p(\bm{z}\mid\bm{\mu}_{1})$ and $p(\bm{z}\mid\bm{\mu}_{2})$, and their cumulative sum approximately recovers the prior distribution. The gray lines connect the samples from the trained models to the corresponding input noise. More details about the toy experiments are given in Appendix D.
@@ 1 Introduction
Generative Models are deep learning algorithms designed to learn the underlying probability distribution of a given dataset, in order to then generate samples coming from such a distribution. Some widely used models are Generative Adversarial Networks (Goodfellow et al., 2014), Variational Autoencoders (VAE; Kingma, 2013; Rezende et al., 2014), and Normalizing Flows (Chen et al., 2018; Papamakarios et al., 2021; Kobyzev et al., 2020). More recently, Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have achieved state-of-the-art (SoTA) results in several domains, including images, videos, and audio (Dhariwal & Nichol, 2021; Rombach et al., 2022; Karras et al., 2022, 2024; Ho et al., 2022; Kong et al., 2021). However, a weakness of diffusion models is the need for an iterative sampling procedure, which can require hundreds of network evaluations. Therefore, substantial effort has been made to develop methods that can maintain similar generation quality while requiring fewer sampling iterations (Song et al., 2021a; Jolicoeur-Martineau et al., 2021; Salimans & Ho, 2022; Liu et al., 2022; Lu et al., 2022). Among such methods, a recent and promising direction is given by Consistency Models (CMs) (Song et al., 2023). CMs, while sharing many similarities with DMs, use a different training procedure as they directly learn the probability flow equations rather than the score function. CMs can be either trained by distilling the ODE trajectories of a pre-trained diffusion model (Consistency Distillation, CD), or completely from scratch through a bootstrap loss (Consistency Training, CT), which results in a novel generative modeling framework. However, the CT objective can be subject to high variance, making it difficult to train. A follow-up work (Song & Dhariwal, 2024) analyses the training dynamics of CMs and proposes several improvements which result in a more stable CT procedure, achieving SoTA results in few-step image generation. Since then, several works have proposed additional strategies to further improve CT (Geng et al., 2024; Wang et al., 2024; Lee et al., 2024b; Yang et al., 2024).
@@
A way to implement this approach is to sample noise using a conditional coupling function between data and noise. The concept of using a coupling function to reduce variance during training was successfully used in Flow Matching, with works such as (Pooladian et al., 2023; Tong et al., 2023; Lee et al., 2023), with the main objective of obtaining straighter ODE trajectories for faster sampling. Forms of coupling in CMs were proposed in works such as (Dou et al., 2024; Issenhuth et al., 2024), but their formulations do not match the performance of standard CMs. In this work, we introduce a variational training of the forward transition kernel with a coupling function between data and noise, which results in a loss function similar to the one used in VAEs. By learning a data-dependent probability distribution over the noise, regularized with an additional Kullback-Leibler divergence loss, we develop an end-to-end training procedure and show how the resulting coupling is effective at improving generation performance and is scalable to high-dimensional data. A simple and intuitive example is shown in Figure 1, where with the same settings, the model trained with the learned coupling generates samples closer to the data distribution. From this figure, we can see how the learned coupling partitions the data differently from how the standard CT does, effectively changing the form of the underlying ODE, likely resulting in an easier training objective as the prior noise is partitioned according to the learned the data-dependent coupling distribution.
@@ 2 Related Work
Since the introduction of Consistency Models in (Song et al., 2023; Song & Dhariwal, 2024), several strategies have been proposed to improve training stability. The work from (Geng et al., 2024) proposes Easy Consistency Models (ECM) a novel training strategy where time steps are sampled in a continuous fashion and the discretization step is adjusted during training, as opposed to the discrete time grid used in iCT. It further shows the benefits of initializing the network weights with the ones from a pretrained score model, achieving superior performance with smaller training budget. (Wang et al., 2024) builds on top of ECM, introducing additional improvements and framing consistency training as value estimation in Temporal Difference learning (Sutton, 2018). Truncated consistency models, introduced in (Lee et al., 2024b), proposes to add a second training stage on top of ECM, to allow the model to focus its capacity on the later time steps, resulting in improved few-steps generation performance. Other recent contributions to the consistency model literature are works such as (Kim et al., 2024b; Heek et al., 2024) where the focus is on improving multistep sample quality, (Lee et al., 2024a) which trains a model with both consistency and score loss to reduce variance, and (Lu & Song, 2024), which proposes several improvements to the continuous-time training of consistency models. Our work can be seen as a parallel contribution to the aforementioned methods, as we focus on learning the data-noise coupling, which can be used as drop-in replacement to the standard independent coupling.
@@
There are several works showing the benefit of using coupling in Flow Matching (Pooladian et al., 2023; Tong et al., 2023; Liu et al., 2023; Lee et al., 2023; Albergo et al., 2024; Kim et al., 2024a). Among these, (Lee et al., 2023) shares the most similarities with our method, as they also use an encoder to learn a probability distribution over the noise conditioned on the data. Their method results in improved performance compared to equivalent Flow Matching models, while requiring less function evaluations. Our method consists of a similar procedure but applied to CT, resulting in improved few-steps generation performance and confirming the effectiveness of learning the data-noise coupling. A different coupling strategy for CT is proposed in (Issenhuth et al., 2024), where the data-noise coupling is extracted directly from the prediction of the consistency model during training. Compared to our method, they do not need the additional encoder to learn the coupling, but their generator-induced coupling needs to be alternated with the standard independent coupling to avoid instabilities. The Flow Matching formulation in Consistency Models with linear interpolation kernel was previously used in (Dou et al., 2024; Yang et al., 2024), where the former also explores the use of minibatch OT coupling, while the latter trains the model to learn the velocity field and adds a regularization term to enforce constant velocity. In our work, we use the Flow Matching formulation, but keep most of the CT building blocks, resulting in a simpler formulation with superior performance.
@@ 3.1 Noise coupling
An advantage of the Flow Matching formalism over SDE diffusion is that, as shown in (Pooladian et al., 2023; Tong et al., 2023), it is straightforward to introduce a probabilistic coupling $\pi(\bm{x}_{1}\mid\bm{x}_{0})$ between the data and the noise distribution. In this case, we require that $\int\pi(\bm{x}_{1}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}$ should follow a standard normal distribution, at least approximately. The use of a non-trivial noise coupling does not alter the form of the conditional velocity fields $\bm{u}_{t}(\bm{x}_{t};\bm{x}_{1})$ as far as the coupling is time-independent. However, it does alter the total velocity field $\bm{u}_{t}(\bm{x}_{t})$ since it affects the conditional distribution $p(\bm{x}_{1},\bm{x}_{t})$ , which determines the expectation in Eq. 2.
@@ 5 Discretized consistency training
6 Consistency Models with learned variational coupling
@@
Here, we restricted our attention to linear forward models of the form $\bm{\psi}_{t}(\bm{x}_{0};\bm{x}_{1})=a_{t}\bm{x}_{0}+b_{t}\bm{x}_{1}$, which encompasses most models used in the diffusion and flow-matching literature. Moreover $\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0})$ and $\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0})$ denote the mean and scale output of the encoder, which define the signal-noise coupling (see Appendix E for a visual representation). Both $\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0})$ and $\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0})$ preserve the same dimensionality of the input signal. The encoder network that produces the coupling distribution $q_{\bm{\phi}}(\bm{x}_{1}\mid\bm{x}_{0})$ can be trained end-to-end alongside the consistency model, as shown in Algorithm 1. This results in a joint optimization where the consistency network adjusts its constancy to minimize its total derivative along the trajectories while the encoder implicitly moves the trajectories towards the space of constancy of the model. In fact, the velocity field $\bm{u}_{t}(\bm{x}_{t})$ depends on the coupling, since $\pi(\bm{x}_{1}\mid\bm{x}_{0})$ affects the expectation in Eq. (2).
Algorithm 1 Variational Consistency Training
@@
This formulation results in a viable generative model as long as the noise at time $1$ remains approximately $p_{1}(\bm{x}_{1})$, since severe deviation from the prior induced by the coupling would result in improper initialization for the sampling procedure and consequently in reduced sample quality. We therefore add a Kullback-Leibler divergence as a regularizer, $\mathcal{D}_{\mathrm{KL}}(q_{\bm{\phi}}||p_{1})$, resulting in a loss resembling the Evidence Lower Bound loss of Variational Autoencoders (Kingma, 2013):
@@
While using an encoder to learn the data-noise coupling requires additional computation during training, we empirically find that a relatively small encoder is enough to learn an effective coupling, which results only in a minor increase of training time (see Appendix C.2). At sampling, the speed and computational requirements are identical to vanilla CMs for the one-step procedure, while for multistep sampling we need to account for additional forward passes of the encoder as shown in Algorithm 2.
@@ 6.1 Connection with variational autoencoders
$\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert^{2}.$
Given that the KL terms in Eq. 7 and the VAE loss are identical, our loss function serves as an upper bound on the loss of a VAE with an encoder $(\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0}),\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0}))$ and a prior $\mathcal{N}(\bm{x}_{1};{\mathbf{0}},{\mathbf{I}})$. Since the VAE loss represents an evidence lower bound, it follows that the consistency loss also provides a lower bound on the model evidence (proof in Appendix B). Compared to traditional VAEs, our method can be viewed as a time-dependent modification where the transition kernel smoothly interpolates between delta distributions centered at datapoints and a Gaussian distribution.
@@ 6.2 Choosing the $\beta$ parameter
$\lambda_{\mathrm{kl}}=\beta\lambda_{\mathrm{ct}}(t_{N})=\frac{1}{\Delta_{t_{N}}}.$
@@ 7 Experiments
In the following, we show that learning the data-noise coupling with our method is a simple yet effective improvement for CT. We pair our Variational Coupling with two established methods as baselines, namely improved Consistency Training (iCT) from (Song & Dhariwal, 2024) and Easy Consistency Tuning (ECM) from (Geng et al., 2024). Note that for the latter, the model is initialized with the weights of a pretrained score model from (Karras et al., 2022), while in the former the weights are initialized at random. More details about the baselines are provided in Appendix C.1. In this work, we consider only the framework of CT, where the unbiased vector field estimator $\bm{u}_{t}(\bm{x}_{t})$ from equation 2 is used to approximate the noisy states $\bm{x}_{t}$ during training, as opposed to Consistency Distillation that uses a pretrained score model as a teacher. We evaluate the models on the image datasets Fashion-MNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), FFHQ $64\times 64$ (Karras, 2019) and (class-conditional) ImageNet $64\times 64$ (Deng et al., 2009). To learn the coupling, we add a smaller version of the neural network used for CT, without time conditioning and with weights always initialized at random. For all the models, we use the variance exploding transition kernel (iCT-VE and ECM-VE) used in (Karras et al., 2022) and (Song & Dhariwal, 2024), with $a_{t}=1$, $b_{t}=t$, and the linear interpolation kernel (iCT-LI and ECM-LI) commonly used in Flow Matching (Lipman et al., 2023), with $a_{t}=1-t/\sigma_{\mathrm{max}}$ and $b_{t}=t/\sigma_{\mathrm{max}}$ (details in Appendix A). For both kernels, we set $\sigma_{\mathrm{min}}=0.002$ and $\sigma_{\mathrm{max}}=80$. More experimental details can be found in Appendix C.2, while samples obtained with our best models are shown in F.

### v2

# VCT: Training Consistency Models with Variational Noise Coupling

@@ Abstract
Consistency Training (CT) has recently emerged as a strong alternative to diffusion models for image generation. However, non-distillation CT often suffers from high variance and instability, motivating ongoing research into its training dynamics. We propose Variational Consistency Training (VCT), a flexible and effective framework compatible with various forward kernels, including those in flow matching. Its key innovation is a learned noise-data coupling scheme inspired by Variational Autoencoders, where a data-dependent encoder models noise emission. This enables VCT to adaptively learn noise-to-data pairings, reducing training variance relative to the fixed, unsorted pairings in classical CT. Experiments on multiple image datasets demonstrate significant improvements: our method surpasses baselines, achieves state-of-the-art FID among non-distillation CT approaches on CIFAR-10, and matches SoTA performance on ImageNet $64\times 64$ with only two sampling steps. Code is available at https://github.com/sony/vct.
@@
Figure 1: Comparison of 1-step generation on toy data for independent and variational coupling. The data is sampled from a 2-d mixture of Gaussians with means $\bm{\mu}_{1}=(0,0.5)$ and $\bm{\mu}_{2}=(0,-0.5)$. On the bottom plot (our approach), we show the posterior probabilities learned by the encoder (in blue and green) corresponding to $p(\bm{z}\mid\bm{\mu}_{1})$ and $p(\bm{z}\mid\bm{\mu}_{2})$, and their cumulative sum approximately recovers the prior distribution. The gray lines connect the samples from the trained models to the corresponding input noise. More details about the toy experiments are given in Appendix E.
@@ 1 Introduction
Generative Models are deep learning algorithms designed to learn the underlying probability distribution of a given dataset, in order to then generate samples coming from such a distribution. Some widely used models are Generative Adversarial Networks (Goodfellow et al., 2014), Variational Autoencoders (VAE; Kingma & Welling, 2013; Rezende et al., 2014), and Normalizing Flows (Chen et al., 2018; Papamakarios et al., 2021; Kobyzev et al., 2020). More recently, Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have achieved state-of-the-art (SoTA) results in several domains, including images, videos, and audio (Dhariwal & Nichol, 2021; Rombach et al., 2022; Karras et al., 2022, 2024; Ho et al., 2022; Kong et al., 2021). However, a weakness of diffusion models is the need for an iterative sampling procedure, which can require hundreds of network evaluations. Therefore, substantial effort has been made to develop methods that can maintain similar generation quality while requiring fewer sampling iterations (Song et al., 2021a; Jolicoeur-Martineau et al., 2021; Salimans & Ho, 2022; Liu et al., 2022; Lu et al., 2022). Among such methods, a recent and promising direction is given by Consistency Models (CMs) (Song et al., 2023). CMs, while sharing many similarities with DMs, use a different training procedure as they directly learn the probability flow equations rather than the score function. CMs can be either trained by distilling the ODE trajectories of a pre-trained diffusion model (Consistency Distillation, CD), or completely from scratch through a bootstrap loss (Consistency Training, CT), which results in a novel generative modeling framework. However, the CT objective can be subject to high variance, making it difficult to train. A follow-up work (Song & Dhariwal, 2024) analyses the training dynamics of CMs and proposes several improvements which result in a more stable CT procedure, achieving SoTA results in few-step image generation. Since then, several works have proposed additional strategies to further improve CT (Geng et al., 2025; Wang et al., 2024; Lee et al., 2025; Yang et al., 2024).
@@
A way to implement this approach is to sample noise using a conditional coupling function between data and noise. The concept of using a coupling function to reduce variance during training was successfully used in Flow Matching, with works such as (Pooladian et al., 2023; Tong et al., 2024; Lee et al., 2023), with the main objective of obtaining straighter ODE trajectories for faster sampling. Forms of coupling in CMs were proposed in works such as (Dou et al., 2024; Issenhuth et al., 2025), but their formulations do not match the performance of standard CMs.
In this work, we propose Variational Consistency Training (VCT), which introduces a variational formulation for training the forward transition kernel via a coupling function between data and noise, a technique we term Variational Coupling (VC). This yields a loss function analogous to that of VAEs. By learning a data-dependent noise distribution regularized with an additional Kullback-Leibler (KL) divergence term, VCT enables an end-to-end training procedure compatible with various forward kernels. We demonstrate that this learned coupling effectively improves generation performance and scales well to high-dimensional data. A simple and intuitive example is shown in Figure 1, where with the same settings, the model trained with the learned coupling generates samples closer to the data distribution. From this figure, we can see how the learned coupling partitions the data differently from how the standard CT does, effectively changing the form of the underlying ODE, likely resulting in an easier training objective as the prior noise is partitioned according to the learned the data-dependent coupling distribution.
@@ 2 Related Work
Since the introduction of Consistency Models in (Song et al., 2023; Song & Dhariwal, 2024), several strategies have been proposed to improve training stability. The work from (Geng et al., 2025) proposes Easy Consistency Models (ECM) a novel training strategy where time steps are sampled in a continuous fashion and the discretization step is adjusted during training, as opposed to the discrete time grid used in iCT. It further shows the benefits of initializing the network weights with the ones from a pretrained score model, achieving superior performance with smaller training budget. (Wang et al., 2024) builds on top of ECM, introducing additional improvements and framing consistency training as value estimation in Temporal Difference learning (Sutton, 2018). Truncated consistency models, introduced in (Lee et al., 2025), proposes to add a second training stage on top of ECM, to allow the model to focus its capacity on the later time steps, resulting in improved few-steps generation performance. Other recent contributions to the consistency model literature are works such as (Kim et al., 2024; Heek et al., 2024) where the focus is on improving multistep sample quality, (Lee et al., 2024) which trains a model with both consistency and score loss to reduce variance, and (Lu & Song, 2025), which proposes several improvements to the continuous-time training of consistency models. Our work can be seen as a parallel contribution to the aforementioned methods, as we focus on learning the data-noise coupling, which can be used as drop-in replacement to the standard independent coupling.
@@
There are several works showing the benefit of using coupling in Flow Matching (Pooladian et al., 2023; Tong et al., 2024; Liu et al., 2023; Lee et al., 2023; Albergo et al., 2024; Kim et al., 2025) and Diffusion Models (Bartosh et al., 2024a, b; Nielsen et al., 2024). Among these, (Lee et al., 2023) shares the most similarities with our method, as they also use an encoder to learn a probability distribution over the noise conditioned on the data. Their method results in improved performance compared to equivalent Flow Matching models, while requiring less function evaluations. Our method consists of a similar procedure but applied to CT, resulting in improved few-steps generation performance and confirming the effectiveness of learning the data-noise coupling. A different coupling strategy for CT is proposed in (Issenhuth et al., 2025), where the data-noise coupling is extracted directly from the prediction of the consistency model during training. Compared to our method, they do not need the additional encoder to learn the coupling, but their generator-induced coupling needs to be alternated with the standard independent coupling to avoid instabilities. The Flow Matching formulation in Consistency Models with linear interpolation kernel was previously used in (Dou et al., 2024; Yang et al., 2024), where the former also explores the use of minibatch OT coupling, while the latter trains the model to learn the velocity field and adds a regularization term to enforce constant velocity. In our work, we use the Flow Matching formulation, but keep most of the CT building blocks, resulting in a simpler formulation with superior performance.
@@ 3.1 Noise coupling
An advantage of the Flow Matching formalism over SDE diffusion is that, as shown in (Pooladian et al., 2023; Tong et al., 2024), it is straightforward to introduce a probabilistic coupling $\pi(\bm{x}_{1}\mid\bm{x}_{0})$ between the data and the noise distribution. In this case, we require that $\int\pi(\bm{x}_{1}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}$ should follow a standard normal distribution, at least approximately. The use of a non-trivial noise coupling does not alter the form of the conditional velocity fields $\bm{u}_{t}(\bm{x}_{t};\bm{x}_{1})$ as far as the coupling is time-independent. However, it does alter the total velocity field $\bm{u}_{t}(\bm{x}_{t})$ since it affects the conditional distribution $p(\bm{x}_{1},\bm{x}_{t})$ , which determines the expectation in Eq. 1.
@@ 5 Discretized consistency training
6 Consistency Models with Variational Coupling
@@
Here, we restricted our attention to linear forward models of the form $\bm{\psi}_{t}(\bm{x}_{0};\bm{x}_{1})=a_{t}\bm{x}_{0}+b_{t}\bm{x}_{1}$, which encompasses most models used in the diffusion and flow-matching literature. Moreover $\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0})$ and $\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0})$ denote the mean and scale output of the encoder, which define the signal-noise coupling (see Appendix F for a visual representation). Both $\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0})$ and $\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0})$ preserve the same dimensionality of the input signal. The encoder network that produces the coupling distribution $q_{\bm{\phi}}(\bm{x}_{1}\mid\bm{x}_{0})$ can be trained end-to-end alongside the consistency model, as shown in Algorithm 1. This results in a joint optimization where the consistency network adjusts its constancy to minimize its total derivative along the trajectories while the encoder implicitly moves the trajectories towards the space of constancy of the model. In fact, the velocity field $\bm{u}_{t}(\bm{x}_{t})$ depends on the coupling, since $\pi(\bm{x}_{1}\mid\bm{x}_{0})$ affects the expectation in Eq. (1).
Algorithm 1 Variational Consistency Training (VCT)
@@
This formulation results in a viable generative model as long as the noise at time $1$ remains approximately $p_{1}(\bm{x}_{1})$, since severe deviation from the prior induced by the coupling would result in improper initialization for the sampling procedure and consequently in reduced sample quality. We therefore add a KL divergence as a regularizer, $\mathcal{D}_{\mathrm{KL}}(q_{\bm{\phi}}||p_{1})$, resulting in a loss resembling the Evidence Lower Bound loss of Variational Autoencoders (Kingma & Welling, 2013):
@@
While using an encoder to learn the data-noise coupling requires additional computation during training, we empirically find that a relatively small encoder is enough to learn an effective coupling, which results only in a minor increase of training time (see Appendix C). At sampling, the speed and computational requirements are identical to vanilla CMs for the one-step procedure, while for multistep sampling we need to account for additional forward passes of the encoder as shown in Algorithm 2.
@@ 6.1 Connection with variational autoencoders
$N\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert^{2}.$
Given that the KL terms in Eq. 7 and the VAE loss are identical, our loss function serves as an upper bound on the loss of a VAE with an encoder $(\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0}),\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0}))$ and a prior $\mathcal{N}(\bm{x}_{1};{\mathbf{0}},{\mathbf{I}})$. Since the VAE loss represents an evidence lower bound, it follows that the consistency loss also provides a lower bound on the model evidence:
Proposition 6.1.
The following upper bound for negative log-density holds:
$-\log p_{\theta}(\bm{x}_{0})$ $\leq\frac{1}{2\sigma^{2}}\,\mathbb{E}_{q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2}+\mathcal{L}_{\mathrm{KL}}(\bm{\phi})$
$\leq\frac{1}{2\sigma^{2}}\int_{0}^{1}\mathbb{E}\left[\left\|\frac{\mathrm{d}}{\mathrm{d}t}\bm{f_{\theta}}(\psi_{t},t)\right\|^{2}\right]\mathrm{d}t+\mathcal{L}_{\mathrm{KL}}(\bm{\phi}),$
where $\mathcal{L}_{\mathrm{KL}}(\bm{\phi}):=\mathcal{D}_{\mathrm{KL}}(q_{\bm{\phi}}(\bm{x}_{1}\mid\bm{x}_{0})||\mathcal{N}(\bm{x}_{1};{\mathbf{0}},{\mathbf{I}}))$.
This proposition also establishes the connection between the minimization objective of CT-VC and that of the continuous-time CM as $N\rightarrow\infty$ (proof in Appendix B).
Compared to traditional VAEs, our method can be viewed as a time-dependent modification where the transition kernel smoothly interpolates between delta distributions centered at datapoints and a Gaussian distribution.
@@ 6.2 Choosing the $\beta$ parameter
$\lambda_{\mathrm{KL}}=\beta\lambda_{\mathrm{ct}}(t_{N})=\frac{\beta}{\Delta_{t_{N}}}.$
@@ 7 Experiments
In the following, we show that learning the data-noise coupling with our method is a simple yet effective improvement for CT. We adapt our Variational Coupling (VC) technique to two established baselines: improved Consistency Training (iCT) (Song & Dhariwal, 2024) and Easy Consistency Tuning (ECM) (Geng et al., 2025). Note that for the latter, the model is initialized with the weights of a pretrained score model from (Karras et al., 2022), while in the former the weights are initialized at random. More details about the baselines are provided in Appendix C.1. In this work, we consider only the framework of CT, where the unbiased vector field estimator $\bm{u}_{t}(\bm{x}_{t})$ from equation 1 is used to approximate the noisy states $\bm{x}_{t}$ during training, as opposed to Consistency Distillation that uses a pretrained score model as a teacher. We evaluate the models on the image datasets Fashion-MNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), FFHQ $64\times 64$ (Karras et al., 2019) and (class-conditional) ImageNet $64\times 64$ (Deng et al., 2009). To learn the coupling, we add a smaller version of the neural network used for CT, without time conditioning and with weights always initialized at random. For all the models, we use the variance exploding transition kernel (iCT-VE and ECM-VE) used in (Karras et al., 2022) and (Song & Dhariwal, 2024), with $a_{t}=1$, $b_{t}=t$, and the linear interpolation kernel (iCT-LI and ECM-LI) commonly used in Flow Matching (Lipman et al., 2023), with $a_{t}=1-t/\sigma_{\mathrm{max}}$ and $b_{t}=t/\sigma_{\mathrm{max}}$ (details in Appendix A). For both kernels, we set $\sigma_{\mathrm{min}}=0.002$ and $\sigma_{\mathrm{max}}=80$. We report more experimental details in Appendix C, comparison with other models in Appendix D, while samples obtained with our best models are shown in Appendix G.